OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 25x2 − 20x − 16 9a2 − 30a − 25 25b2 − 15b + 9 16x2 − 24x + 9

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

explain please?

OpenStudy (angela210793):

Jimmy's right(as always) 16x^2-24x+9=(4x-3)^2

OpenStudy (anonymous):

16x2 − 24x + 9 = (4x -3)(4x-3)

OpenStudy (anonymous):

ty

OpenStudy (dumbcow):

(a^2 -2ab +b^2) = (a-b)^2 a^2 and b^2 must be perfect squares middle term must be double the product of the roots a,b

OpenStudy (anonymous):

oh angela! - you cant have seen all my answers!!

OpenStudy (angela210793):

well...ur right at least 99% of the answers i've seen :)